ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 62 Kim, L. V., Institute of Mathematics of the Uzbek Academy, Uzbekistan Generalized model of random censorship from both sides A generalised model of random censorship from the left and right is introduced and some information properties of statistics constructed by the data, censored according to the model, are studied. Cramer-Rao type lower bound is considered and the conditions of attaining of the bound are given under the presence of nuisance parameters. The model of proportional hazard for the scheme of censoring is considered as well. Reference: L. Horvath, On random censorship from both sides, Math. Operation. u. Stat. ser. Stat. 15(1984)4, 581-594.